I can get Computer Eng ,EEE&ECE in Guwahati and CSE &Electrical Engineering in IIT Varnasi . Which one is preferable?

Grade:12

2 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
I think ECE and EEE are good branches to take in IIT guwahati but if you are inclined toward CS and want to get high salary jobs you should join CS in Varansi.
But being in old IITs is a different thing too.
Arun Kumar
Btech
IIT delhi
Askiitians Faculty
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Phenol dimerises in the solvent (organic) as :
2C6H5OH ⇌ (C6H5OH)2
1 mol 0 mol (Initial)
1 - ∝ mol ∝/2 mol (Equilibrium)
Van’t Hoff factor, i = 1 - ∝ + ∝/2 = 1 - ∝/2
Where ∝ is the degree of dimerization.
∆Tf(obs) = Kf * molality * i
Or 7 = 14 * m * (1 - ∝/2)
In the given problem, amount of the solvent used to dissolved 75.2 g of phenol is not mentioned. Hence molality m, can not be calculated and so also the value of ∝.
If amount of solution is presumed to be 1 kg (= 100g) then,
M = 75.2/94 = 0.8 (mol. mass of phenol = 9.4)
And 7 = 14 * 75.2/94 * (1 - ∝/2)
Or ∝ = 0.75 = 75%

Thanks and Regards